# Conceptualization of this study, Methodology, Writing

### v1

@@ Abstract
Understanding actions within surgical workflows is essential for evaluating post-operative outcomes. However, capturing long sequences of actions performed in surgical settings poses challenges, as individual surgeons have their unique approaches shaped by their expertise, leading to significant variability. To tackle this complex problem, we focused on segmentation with precise boundaries, a demanding task due to the inherent variability in action durations and the subtle transitions often observed in untrimmed videos. These transitions, marked by ambiguous starting and ending points, complicate the segmentation process. Traditional models, such as MS-TCN, which depend on large receptive fields, frequently face challenges of over-segmentation (resulting in fragmented segments) or under-segmentation (merging distinct actions). Both of these issues negatively impact the quality of segmentation. To overcome these challenges, we present the Multi-Stage Boundary-Aware Transformer Network (MSBATN) with hierarchical sliding window attention, designed to enhance action segmentation. Our proposed approach incorporates a novel unified loss function that treats action classification and boundary detection as distinct yet interdependent tasks. Unlike traditional binary boundary detection methods, our boundary voting mechanism accurately identifies start and end points by leveraging contextual information. Extensive experiments using three challenging surgical datasets demonstrate the superior performance of the proposed method, achieving state-of-the-art results in F1 scores at thresholds of 25% and 50%, while also delivering comparable performance in other metrics.
@@ 1 Introduction
Surgical action segmentation involves analyzing complex elements such as instruments, organs, fluids, or smoke to identify and categorize distinct actions or phases in surgical videos Mascagni et al. (2022). This task is crucial for enhancing surgical training by providing detailed feedback to trainees, improving workflow efficiency and patient safety through real-time monitoring and optimization of procedures, and advancing the performance of autonomous or semi-autonomous surgical robots capable of understanding and executing complex tasks. Recent advancements in this domain have led to the creation of sophisticated architectures tailored to the unique challenges of surgical environments (like Gazebo, ROS2 software development kit (SDK), etc), enabling better decision-making services during operations and driving progress in surgical robotics and education Kiyasseh et al. (2023).
Surgical videos are often heavily annotated manually, where each frame is labelled to a surgical phase or sub-phase or specific action/gesture, as well as the surgical instruments present in that frame Psychogyios et al. (2023); Twinanda et al. (2016). Recent datasets have incorporated action triplets Nwoye et al. (2022), which include action, instrument, and organ information for frame-wise action recognition and instruments or organs segmentation Nwoye et al. (2023); Chadebecq et al. (2023). However, temporal evolution hierarchies are not well defined as they are subjective in nature. The high to low-level temporal component of a surgical operation or intraoperative activity is termed phase, steps and tasks. For more details please refer attached supplementary material file.
Figure 1 demonstrates an effective approach to handle temporal variations by modelling start-end transition in long videos. It starts with frame-wise action and boundary recognition, and then it goes through a segment-wise robust refinement process to avoid over-under segmentation.
For efficient analysis of surgical workflows and improving postoperative outcomes, a robust understanding of both spatial context and temporal relationships, the arrangement and transitions between actions or gestures, are essential Chadebecq et al. (2023). In this regard, most publicly available datasets often lack these temporal hierarchies and their evolutions. Additionally, similarities between different phases, including unclear or ambiguous starting, ending and transitions between actions 2, lead to inaccurate recognition.
Figure 1: Overview of the proposed framework, which focuses on segments encompassing the action’s start, centre, and end, enabling the model to capture nuanced transitions and improve boundary precision.
For improved surgical video analysis, most existing models are built based on deep neural networks which are broadly divided into single-stage and multi-stage where the architectures evolved from RNN to Temporal Convolution Network(TCN)Farha and Gall (2019); Li et al. (2020); Czempiel et al. (2020); Yi et al. (2022), and Transformer Yang et al. (2024); Zhang et al. (2022); Ding and Li (2022); Nwoye et al. (2022); Shi et al. (2023) based approaches. Single-stage models make predictions directly from visual-temporal features. These models utilize a small window of time around the target frame to capture spatial-temporal features, which minimizes computational costs. Early models like EndoNetTwinanda et al. (2016); Twinanda (2017) focused on detecting tool presence and recognizing phases using RNN-based methods. EndoNetTwinanda et al. (2016) is the first method proposed for two tasks simultaneously in a multitasking manner, including tool presence detection and phase recognition. EndoLSTM Twinanda (2017), as the improved EndoNet, replaced hierarchical hidden Markov models (HHMMs) with LSTM for temporal refinement and improved the results substantially. Besides, in Mondal et al. (2019), a multitask deep learning framework was proposed comprising ResNet-50 and Bi-LSTM. It performs better in tool detection and phase recognition. In Jin et al. (2020), a single-staged multitask recurrent convolutional network was proposed with correlation loss (MTRCNet-CL) for surgical video analysis. This multitask learning method used a new correlation loss to minimize tool and phase prediction divergences.
However, RNN and LSTM-based approaches often struggle to capture long-term dependencies that are essential for accurately modelling the entire surgical process. For instance, models like Actionformer Zhang et al. (2022) and Tridet Shi et al. (2023) integrate Transformer architecture for action recognition. However, these single-stage models are designed for small actions while each phase or action segment in surgery takes several minutes, or the whole procedure sometimes takes more than 1 hour. Because of this research, the signal-stage models are not ideal for long untrimmed video analysis.
Multi-stage models, on the other hand, include an additional refinement stage to improve the initial predictions by correcting the errors. Multi-stage models are considered more suitable for surgical phase recognition to optimize the complexity and ordered nature of surgical activitiesCzempiel et al. (2020); Van Amsterdam et al. (2022). Multi-stage models such as TeCNOCzempiel et al. (2020) and Farha and Gall (2019); Li et al. (2020) employ multiple temporal convolutional networks to refine predictions in subsequent stages. However, these methods often neglect high-level semantic information and struggle to refine predictions effectively in an end-to-end manner, leading to over or under-segmentation errors. This is because the refinement stage focuses on adjusting initial predictions without fully understanding the actions involved during training, and errors remain during inference. This can lead to misinterpretations in complex scenarios where context is crucial. Recently, in Ramesh et al. (2021), a multitask multistage temporal convolutional network (MTMS-TCN) was proposed to jointly learn the tasks of phase and step recognition. Some studies applied sensor data with attention mechanism to enhance the understanding of surgical processes Van Amsterdam et al. (2022); Goldbraikh et al. (2023), however aligning sensor(kinematic) data with corresponding video frames is complex due to differences in sampling rates and potential time lags(different frequencies), leading to misalignment issues. To improve the segmentation, some research has focused on incorporating segment-level modeling with transformer Ding and Li (2022); Jin et al. (2022). Their main idea is to use high-level segments from surgical videos to clarify ambiguous semantics presented in neighboring frames.
Figure 2: Ambiguous action ordering in untrimmed long surgical videos
@@
To overcome the above limitations, the current methods use sliding window detectors Czempiel et al. (2020), but they struggle to capture long-term patterns effectively Twinanda et al. (2016); Twinanda (2017); Mondal et al. (2019); Czempiel et al. (2020). Additionally, these approaches overlook the finer details that are essential for understanding the intricate relationships between spatial and temporal features Kiyasseh et al. (2023), resulting in over and under-segmentation errors. The over-segmentation occurs when a segment is incorrectly divided into multiple action segments, leading to the detection of more actions than actually exist. These approaches are overly sensitive to minor variations in the input data, causing them to interpret slight changes as new action boundaries Ishikawa et al. (2021). On the other hand, Under-segmentation happens when distinct actions are mistakenly merged into a single segment. This typically arises due to a lack of sufficient temporal resolution or failure to recognize subtle transitions between different actions. So, the model overlooks actual action boundaries, resulting in fewer detected action segments than actually exist.
@@
Most of the approaches fail to model the continuous and smooth transitions characteristic of surgical activities due to treating actions and boundaries as discrete events without accounting for their gradual shifts between actions Fard et al. (2017); Ishikawa et al. (2021). We noticed that the self-attention mechanism has difficulty focusing on important local areas, particularly at the initial layers. Additionally, the encoder-decoder architectures do not effectively refine predictions based on the temporal relationships between actions Yi et al. (2021). To address these issues, first, we propose a framework that extracts spatial features with DinoV2 by discarding irrelevant large patch tokens with register tokens Oquab et al. (2023). Second, for temporal modelling, we present an encoder-decoder model that is built based on the combination of dilated TCN, unified hierarchical and sliding window attention, which captures coarse-to-fine context. The hierarchical representation with dynamic sliding window attention is used to handle long input sequences that organize the attention layers hierarchically to first focus on local relationships and then gradually capture global dependencies. Last, we develop a unified loss function by combining the loss functions of individual tasks into a single action segmentation to achieve effective boundary detection. This holistic approach allows one to learn multiple tasks simultaneously, leveraging shared representations to improve overall segmentation performance.
Our approach introduces the following key innovations:
@@
  • We introduce hierarchical sliding window attention mechanisms (HSWA) to resolve ambiguities among action orders from different temporal resolution. It links high-level action segment structures with fine-grained frame-wise details, enabling the model to better understand long-term temporal dependencies and immediate frame-level changes.
  • We propose a boundary aware loss function that gradually measures cosine similarity from center to boundaries inside action segments that assist in minimizing over or under-segmentation error.
@@ 2.1 Action Recognition and Segmentation
For the surgical domain, EndoNetTwinanda et al. (2016) was developed for phase recognition in cholecystectomy videos by learning inherent visual features through a CNN architecture. Recognizing the importance of tool information in phase recognition, EndoNet adopted a multi-task framework that incorporates tool presence detection into the feature learning process. An extension of AlexNet was designed with dual objectives: one part focused on detecting the presence of surgical tools, while the other was responsible for recognizing seven distinct surgical phases. To capture temporal dependencies, EndoNet integrated a Hierarchical Hidden Markov Model (HMM), enabling the modelling of temporal transitions between phases. The architecture is further improved by LSTMTwinanda (2017) and ResnetTwinanda et al. (2018). However, these methods struggle with accurately identifying transition frames between action phases. To overcome this limitation, a sequence-to-sequence model with an attention mechanism was proposed Namazi et al. (2019); Yi et al. (2021), which accommodates variable-length inputs and outputs, making it well-suited for detecting overlapping or sequential surgical tasks.
Temporal Action Detection (TAD) methodologies are mainly classified into two-stage and one-stage-based methods. The two-stage approach, inspired by object detection frameworks like Faster R-CNN, involves an initial phase of generating action proposals, followed by subsequent stages of classification and refinement Chao et al. (2018); Keisham et al. (2022). Although this method often achieves higher accuracy due to its detailed localization process, it tends to be computationally intensive, posing challenges for real-time applications Yi et al. (2022). On the other hand, the one-stage approach together consolidates action proposal generation and classification in a single framework, which offers computational efficiency for real-time scenarios. For instance, models like ActionFormer Zhang et al. (2022), and Surgformer Yang et al. (2024) are transformer-based architectures that have achieved enhanced precision of action localization while maintaining computational efficiency. These models capture long-range dependencies and contextual information that helps improve the accuracy of action segmentation and detection. However, one-stage methods have historically faced difficulties in precisely localizing action boundaries in long untrimmed videos due to their small receptive fields, which is crucial for optimal performance in TAD.
@@ 2.2 Boundary Detection
Identifying action boundaries within untrimmed videos poses unique challenges due to the inherent ambiguity in determining precise start and end points of actions. Initial action boundary detection approaches used probabilistic models to capture temporal dynamics. For example, a spatiotemporal Gaussian Mixture Model (GMM) with Variational Bayesian (VB) Loukas et al. (2016) inference was employed to dynamically estimate the number of components in the GMM, which eliminates the need to predefine the number of clusters. VB inference approximated the full posterior distribution, which improved model flexibility and reduced redundant components. Label assignment and tracking were achieved through Kullback-Leibler (KL) distance, ensuring consistent tracking across video clips. In contrast to rigid boundary detection methods, the Soft-Boundary Unsupervised Gesture Segmentation (Soft-UGS) Fard et al. (2017) approach was proposed as a soft-boundary clustering technique that allowed for gradual transitions between actions. This method utilized kinematic trajectory data, Dynamic Time Warping (DTW), and probabilistic clustering to segment surgical gestures effectively in real-time applications. While these probabilistic models laid the groundwork for more advanced methods, they had limitations in precision, particularly in detecting abrupt transitions in video content.
@@
On the other hand, the advent of deep learning introduced new techniques for addressing the temporal and spatial complexities of action boundary detection. For the surgical domain, sequence models were applied to map variable-length input sequences to output sequences while incorporating attention mechanisms for more accurate boundary localization. A double attention mechanism(such as Namazi et al. (2019)) was introduced to focus specifically on the beginning and end of each surgical phase, offering a refined approach to boundary detection in complex tasks like surgery.
@@
The above methods for temporal segmentation face the issue of over-segmentation because they segment actions into too many sub-components, leading to inaccurate predictions. The Action Segment Refinement Framework (ASRF)Ishikawa et al. (2021) was introduced to address this problem by decoupling frame-wise classification from boundary regression. ASRF refines action segmentation through two branches: one for action predictions and another for boundary regression. By focusing specifically on predicted action boundaries, ASRF achieved overall action localization accuracy. ASRF*Yi et al. (2021) improved the performance with an attention mechanism.
@@
Recent methods, such as TriDetShi et al. (2023) and Elastic Moment Bounding (EMB)Huang et al. (2022), have further extended the capabilities of one-stage detection for real-time applications. TriDet introduced the Trident-head detection mechanism, which models action boundaries by estimating relative probability distributions instead of directly predicting offsets. This innovation significantly improved boundary localization accuracy while reducing computational overhead. EMB tackles uncertainties in temporal boundaries by transforming rigid and manually labelled endpoints into an elastic set of boundaries; however, it does tend to generate a high number of false positives.
@@
Building upon these advancements, our action boundary network employs a voting mechanism to improve segmentation by accurately finding the start and end of an action. By utilizing the predicted action boundaries, our framework refines frame-wise predictions, ultimately improving the performance of action segmentation.
Table 1: Notation table
Symbol Description
$T$ Sequence length (number of frames).
$D$ Dimensionality of frame features.
$\mathbf{x}_{t}$ Feature vector of the $t$-th frame.
$C$ Set of action categories (including background).
$S$ Set of action segments.
$W$ Base window size for sliding attention.
$d$ Dilation factor for dilated attention.
$T_{s}$ Sequence length at scale $s$.
$\alpha_{ij}^{s}$
Attention weight for token $x_{i}$ attending
to $x_{j}$ at scale $T_{s}$.
$\mathcal{L}_{\text{dice}}$ Dice loss for segment overlap.
$\mathcal{L}_{\text{sim}}$ Gaussian-weighted cosine similarity loss.
$\mathcal{L}_{\text{boundary}}$ Boundary-aware loss for segment precision.
$\alpha,\beta,\gamma,\delta$ Weights for different loss components.
$G(t,t+1)$ Gaussian weight emphasizing temporal proximity.
$G_{\text{boundary}}(t)$ Gaussian weight centered on segment boundaries.

### v2

@@ Abstract
Understanding actions within surgical workflows is critical for evaluating post-operative outcomes and enhancing surgical training and efficiency. Capturing and analyzing long sequences of actions in surgical settings is challenging due to the inherent variability in individual surgeon approaches, which are shaped by their expertise and preferences. This variability complicates the identification and segmentation of distinct actions with ambiguous boundary start and end points. The traditional models, such as MS-TCN, which rely on large receptive fields, that causes over-segmentation, or under-segmentation, where distinct actions are incorrectly aligned. To address these challenges, we propose the Multi-Stage Boundary-Aware Transformer Network (MSBATN) with hierarchical sliding window attention to improve action segmentation. Our approach effectively manages the complexity of varying action durations and subtle transitions by accurately identifying start and end action boundaries in untrimmed surgical videos. MSBATN introduces a novel unified loss function that optimises action classification and boundary detection as interconnected tasks. Unlike conventional binary boundary detection methods, our innovative boundary weighing mechanism leverages contextual information to precisely identify action boundaries. Extensive experiments on three challenging surgical datasets demonstrate that MSBATN achieves state-of-the-art performance, with superior F1 scores at 25% and 50% thresholds and competitive results across other metrics.
@@ 1 Introduction
Surgical action segmentation involves analyzing instruments, organs, fluids, and smoke to identify and categorise distinct actions or phases in complex surgical videos Mascagni et al. (2022). This process is essential to improve surgical training by providing detailed feedback to trainees and advancing the performance of autonomous surgical robots. These robots can understand and perform intricate tasks, thus improving workflow efficiency and patient safety through real-time monitoring and optimisation of procedures. Recent advances in this field have led to the development of sophisticated architectures tailored to the unique challenges of surgical environments, such as Gazebo and the ROS2 software development kit (SDK). These innovations support better decision-making during operations and drive progress in surgical robotics and education Kiyasseh et al. (2023).
Figure 1: The action-transition graph for SAR-RARP50 features nodes 1 to 7, representing canonical suturing actions. A typical sequence is: Pick up the needle (1) → Position the tip (2) → Push through the tissue (3) → Pull out (4). After this sequence, the surgeon can either return or drop the needle (7), tie a knot (5), cut the suture (6), or skip the last two steps. This graph allows for loops and re-entries, so it doesn’t follow a single linear path from start to finish.
Surgical videos are typically annotated manually, with each frame labeled according to the surgical phase, sub-phase, or specific actions/gestures, as well as the instruments present in that frame Psychogyios et al. (2023); Twinanda et al. (2016). Recent datasets have introduced action triplets Nwoye et al. (2022), which include information about the action, instruments, and organs for frame-wise action recognition and segmentation of instruments or organs Nwoye et al. (2023); Chadebecq et al. (2023). Most of these datasets focus primarily on spatial information, while providing minimal temporal information.
There is a standard temporal hierarchy that surgeons follow during procedures. The highest-level temporal component of a surgical operation or intraoperative activity is referred to as a "phase," which encompasses a series of related activities. Each phase consists of several "steps," which are the specific actions taken to achieve the goals of that phase. Each step represents a procedure-specific segment aimed at addressing the patient’s condition, while the "task" is a sub-component of a step that involves a series of actions Meireles et al. (2021). However, the definitions of temporal evolution hierarchies can be subjective and depend on the surgeon’s experience, which is why they are rarely disclosed alongside the videos. This lack of structure, coupled with visual similarities between actions and fuzzy boundaries, often leads to ambiguous transitions (Figure 1). For example, in SAR-RARP50, the action-transition graph is highly cyclic: any of the seven suturing gestures can be repeated or skipped, resulting in frequent re-entries and out-of-order sequences.
Deep learning pipelines for surgical video analysis have advanced significantly, evolving from recurrent networks to temporal convolutions, and more recently, to Transformers Farha and Gall (2019); Li et al. (2020); Czempiel et al. (2020); Yi et al. (2022); Yang et al. (2024); Zhang et al. (2022); Ding and Li (2022); Nwoye et al. (2022); Shi et al. (2023). Early systems, such as EndoNet Twinanda et al. (2016) and EndoLSTM Twinanda (2017), utilised CNN backbones for frame features and RNN modules for temporal smoothing, achieving the first joint modeling of tool presence and phase recognition. Subsequent variants incorporated correlation losses, bidirectional LSTMs, or multitask heads to improve accuracy Mondal et al. (2019); Jin et al. (2020). However, RNN-based designs often struggle to retain dependencies over minutes—to hours—long procedure.
To expand the temporal receptive field without encountering the vanishing gradient problems associated with RNNs, researchers have shifted to Temporal Convolutional Networks (TCNs)) Farha and Gall (2019); Li et al. (2020); Czempiel et al. (2020); Yi et al. (2022). The MS-TCN family employs dilated 1-D kernels that efficiently aggregate historical information while maintaining a low inference latency Farha and Gall (2019); Li et al. (2020). Recently, single-stage Transformers like Actionformer Zhang et al. (2022) and TriDet Shi et al. (2023) have provided global self-attention over short clips, excelling in fine-grained gesture localisation. However, these single-pass models were designed for brief human actions; their fixed window makes them unsuitable to capture the hierarchical nature of surgical workflows. Multi-stage architectures address this limitation by incorporating one or more refinement passes to correct over- or under-segmentation produced in the first stage. TeCNO Czempiel et al. (2020) chains several TCN blocks, each trained to refine the output of its predecessor, while MTMS-TCN jointly learns phases and steps to leverage shared cues Farha and Gall (2019); Li et al. (2020); Ramesh et al. (2021). Other research has integrated kinematic sensor streams or attention mechanisms to improve temporal reasoning Van Amsterdam et al. (2022); Goldbraikh et al. (2023), and segment-level Transformers have refined ambiguous boundaries by analysing high-level proposals Ding and Li (2022); Jin et al. (2022). Despite these advancements, refinement modules often work with hard predictions without an explicit model of surgical semantics, which can lead to persistent systematic errors, particularly when actions visually resemble one another.
Figure 2: Multi-stage Boundary-Aware Refinement Pipeline: The initial prediction with typical over- and under-segmentation errors (segments are split too finely, merged, or slightly shifted) misleads action order. Our approach resolves this by applying a Gaussian centre-weighting scheme to Action and Boundary segments to minimise the error. The refined prediction aligns with the annotated ground truth, and boundaries without re-processing the video.
@@
Furthermore, these approaches overlook the finer details that are essential for understanding the intricate relationships between spatial and temporal features Kiyasseh et al. (2023), resulting in over and under-segmentation errors. The over-segmentation occurs when a segment is incorrectly divided into multiple action segments, leading to the detection of more actions than actually exist. These approaches are overly sensitive to minor variations in the input data, leading them to interpret slight changes as new action boundaries Ishikawa et al. (2021). On the other hand, under-segmentation happens when distinct actions are mistakenly merged into a single segment. This typically arises due to a lack of sufficient temporal resolution or the inability to recognise subtle transitions between different actions. So, the model overlooks the actual action boundaries, resulting in fewer detected action segments than actually exist.
The continuous and smooth transitions characteristic of surgical activities are often ignored due to treating actions and boundaries as discrete events without accounting for their gradual shifts between actions Fard et al. (2017); Ishikawa et al. (2021). We noticed that the self-attention mechanism has difficulty focusing on important local areas, particularly at the initial layers. Additionally, the encoder-decoder architectures do not effectively refine predictions based on the temporal relationships between actions Yi et al. (2021). Our solution (Fig. 2) tackles these problems in two steps: firstly, a backbone network predicts initial labels for actions and their boundaries. Secondly, a lightweight post-processing module corrects over- and under-segmentation by focusing exclusively on the start–end transitions. Because it operates only on predicted segments, it adds negligible computational cost to the main model. Training utilises a duration-aware loss weighting that adjusts each segment’s contribution inversely based on its duration, ensuring that short yet clinically significant actions are not overshadowed by longer ones.
@@
We propose a framework that extracts spatial features with DinoV2 by discarding irrelevant large patch tokens with register tokens Oquab et al. (2023). For temporal modelling, we present an encoder-decoder architecture that is built based on the combination of dilated TCN, unified hierarchical and sliding window attention, which captures fine-to-coarse context. The hierarchical representation with dynamic sliding window attention is used to handle long input sequences that organise the attention layers hierarchically to first focus on local relationships and then gradually capture global dependencies. We design a unified loss function by combining the loss functions of individual tasks into a single action segmentation to achieve effective boundary detection. Our approach introduces the following key innovations:
@@
  • We introduce hierarchical sliding window attention mechanisms (HSWA) to resolve ambiguities among action orders from different temporal resolutions. It links high-level action segment structures with fine-grained frame-wise details, enabling the model to better understand long-term temporal dependencies and immediate frame-level changes.
  • We propose a boundary-aware loss function that gradually measures cosine similarity from centre to boundaries inside action segments that assist in minimising over- or under-segmentation error.
@@ 2.1 Action Recognition and Segmentation
For the surgical domain, EndoNet Twinanda et al. (2016) was developed for phase recognition in cholecystectomy videos by learning inherent visual features through a CNN architecture. Recognising the importance of tool information in phase recognition, EndoNet adopted a multi-task framework that incorporates tool presence detection into the feature learning process. An extension of AlexNet was designed with dual objectives: one part focused on detecting the presence of surgical tools, while the other was responsible for recognising seven distinct surgical phases. To capture temporal dependencies, EndoNet integrated a Hierarchical Hidden Markov Model (HMM), enabling the modelling of temporal transitions between phases. The architecture is further improved by LSTM Twinanda (2017) and ResNet Twinanda et al. (2018). However, these methods struggle with accurately identifying transition frames between action phases. To overcome this limitation, a sequence-to-sequence model with an attention mechanism was proposed Namazi et al. (2019); Yi et al. (2021), which accommodates variable-length inputs and outputs, making it well-suited for detecting overlapping surgical tasks.
Temporal Action Detection (TAD) methodologies are mainly classified into two-stage and one-stage-based methods. The two-stage approach, inspired by object detection frameworks like Faster R-CNN, involves an initial phase of generating action proposals, followed by subsequent stages of classification and refinement Chao et al. (2018); Keisham et al. (2022). Although this method often achieves higher accuracy due to its detailed localisation process, it tends to be computationally intensive, posing challenges for real-time applications Yi et al. (2022). On the other hand, the one-stage approach consolidates action proposal generation and classification in a single framework, which offers computational efficiency for real-time scenarios. Models like ActionFormer Zhang et al. (2022), and Surgformer Yang et al. (2024) are transformer-based architectures that have achieved enhanced precision of action localisation while maintaining computational efficiency. These models capture long-range dependencies and contextual information that helps improve the accuracy of action segmentation and detection. However, one-stage methods have historically faced difficulties in precisely localising action boundaries in long untrimmed videos due to their small receptive fields, which is crucial for optimal performance in TAD.
@@ 2.2 Boundary Detection
Identifying action boundaries within untrimmed videos poses unique challenges due to the inherent ambiguity in determining precise start and end points of actions. Initial action boundary detection approaches used probabilistic models to capture temporal dynamics. For example, a spatiotemporal Gaussian Mixture Model (GMM) with Variational Bayesian (VB) Loukas et al. (2016) inference was employed to dynamically estimate the number of components in the GMM, which eliminates the need to predefine the number of clusters. VB inference approximated the full posterior distribution, which improved model flexibility and reduced redundant components. Label assignment and tracking were accomplished using Kullback-Leibler (KL) distance, ensuring consistent tracking across video clips. Unlike rigid boundary detection methods, the Soft-Boundary Unsupervised Gesture Segmentation (Soft-UGS) approach Fard et al. (2017) was proposed as a soft-boundary clustering technique that allowed for gradual transitions between actions. This method effectively utilised kinematic trajectory data, Dynamic Time Warping (DTW), and probabilistic clustering to segment surgical gestures in real-time applications. Although these probabilistic models provided a foundation for more advanced techniques, they faced limitations in precision, especially in detecting abrupt transitions in video content.
@@
In contrast, the emergence of deep learning introduced new techniques for addressing the temporal and spatial complexities of action boundary detection. For the surgical domain, sequence models were applied to map variable-length input sequences to output sequences while incorporating attention mechanisms for more accurate boundary localisation. A double attention mechanism(such as Namazi et al. (2019)) was introduced to focus specifically on the beginning and end of each surgical phase, offering a refined approach to boundary detection in complex tasks like surgery.
@@
The above methods for temporal segmentation face the issue of over-segmentation because they segment actions into too many sub-components, leading to inaccurate predictions. The Action Segment Refinement Framework (ASRF) Ishikawa et al. (2021) was introduced to address this problem by decoupling frame-wise classification from boundary regression. ASRF refines action segmentation through two branches: one for action predictions and another for boundary regression. By focusing specifically on predicted action boundaries, ASRF achieved overall action localisation accuracy. ASRF* Yi et al. (2021) improved the performance with an attention mechanism.
@@
Recent methods, such as TriDet Shi et al. (2023) and Elastic Moment Bounding (EMB) Huang et al. (2022), have further extended the capabilities of one-stage detection for real-time applications. TriDet introduced the Trident-head detection mechanism, which models action boundaries by estimating relative probability distributions instead of directly predicting offsets. This innovation significantly improved boundary localisation accuracy while reducing computational overhead. EMB tackles uncertainties in temporal boundaries by transforming rigid and manually labelled endpoints into an elastic set of boundaries; however, it does tend to generate a high number of false positives.
@@
Building upon these advancements, our action boundary network employs a voting mechanism to improve segmentation by accurately finding the start and end of an action. By utilising the predicted action boundaries, our framework refines frame-wise predictions, ultimately improving the performance of action segmentation.